# The Open mulTiwavelength Transient Event Repository (OTTER): Infrastructure Release and Tidal Disruption Event Catalog

### v1

@@ Appendix A Metadata and References for each TDE
[274] Jones et al. (2021) [275] Munoz-Arancibia et al. (2021a) [276] Munoz-Arancibia et al. (2021g)
[277] Munoz-Arancibia et al. (2021h) [278] Fremling (2021a) [279] Forster et al. (2021b)
[280] Fremling (2021b) [281] Hodgkin et al. (2021) [282] Wevers et al. (2022)
[283] Horesh et al. (2022) [284] Liu et al. (2022) [285] Perlman et al. (2022)
[286] Ravi et al. (2022) [287] Mohan et al. (2022) [288] Sfaradi et al. (2022)
[289] Yao et al. (2022f) [290] Cendes et al. (2022) [291] Wang et al. (2022)
[292] Goodwin et al. (2022) [293] Angus et al. (2022) [294] Andreoni et al. (2022)
[295] Yao et al. (2022a) [296] Fulton et al. (2022b) [297] Chu et al. (2022)
[298] Tucker (2022) [299] Arcavi et al. (2022b) [300] Arcavi et al. (2022a)
[301] Fulton et al. (2022a) [302] Hinkle (2022) [303] Yao et al. (2022c)
[304] Kangas (2022) [305] Moore et al. (2022) [306] Hammerstein et al. (2022a)
[307] Yao et al. (2022e) [308] Somalwar et al. (2022) [309] Yao et al. (2022b)
[310] Blanchard et al. (2022) [311] Hammerstein et al. (2022b) [312] Fulton et al. (2022c)
[313] Yao (2022a) [314] Yao et al. (2022d) [315] Yao (2022b)
[316] Fulton et al. (2022d) [317] Tonry et al. (2022a) [318] Munoz-Arancibia et al. (2022a)
[319] Hodgkin et al. (2022) [320] Tonry et al. (2022b) [321] Tonry et al. (2022c)
[322] De (2022) [323] Andreoni (2022) [324] Tonry et al. (2022d)
@@
[328] Fremling (2022a) [329] Munoz-Arancibia et al. (2022c) [330] Stanek (2022)
[331] Tonry et al. (2022f) [332] Tonry et al. (2022g) [333] Fremling (2022b)
[334] Fremling (2022c) [335] Homan et al. (2023) [336] Hammerstein et al. (2023a)
[337] Panagiotou et al. (2023) [338] Zhu et al. (2023) [339] Yao et al. (2023)
[340] Hammerstein et al. (2023b) [341] Aleo et al. (2023) [342] Almeida et al. (2023)
[343] Goodwin et al. (2023a) [344] Rhodes et al. (2023) [345] Goodwin et al. (2023b)
[346] Goldtooth et al. (2023) [347] Yao et al. (2023a) [348] Wise et al. (2023)
[349] Chornock et al. (2023) [350] Hammerstein et al. (2023a) [351] Taguchi et al. (2023b)
[352] Yao et al. (2023b) [353] Taguchi et al. (2023a) [354] Charalampopoulos et al. (2023)
@@
[358] Yao (2023) [359] Chornock (2023) [360] Hammerstein et al. (2023b)
[361] Stanek (2023) [362] Tonry et al. (2023) [363] Fremling (2023)
[364] Sollerman (2023) [365] Munoz-Arancibia et al. (2023) [366] Somalwar et al. (2025)
[367] Charalampopoulos et al. (2024) [368] Masterson et al. (2024) [369] Zhang et al. (2024)
[370] Wevers et al. (2024) [371] Guolo et al. (2024) [372] Faris et al. (2024)
[373] Cendes et al. (2024) [374] Dykaar et al. (2024) [375] Christy et al. (2024)
[376] Anumarlapudi et al. (2024) [377] Yao et al. (2024) [378] Bykov et al. (2024)
@@
[385] Sollerman et al. (2024) [386] Somalwar et al. (2025) [387] Hajela et al. (2025)
[…]
@@ Appendix B Detailed Data Schema
    classification: All reported classifications of this transient stored as a list. Every classification has a confidence associated with it which is derived analytically based on the information available in OTTER. This is described in more detail in §2.4.
@@ B.6 Properties
The classification property is a dictionary with meta-level flag keywords and then all of the classifications stored under the value keyword. The current meta-level flags are spec_classed, unambiguous, and discovery_method, and are discussed in detail in §2.4. In addition to the keywords available in §B.3, classification value elements elements can have: object_class, confidence (which is a flag, as defined in §2.4), and ml_score. An object can have multiple classifications.
@@
Appendix C Non-cumulative TDE Discovery Histogram
Figure 16 shows a non-cumulative version of Figure 7. This is a different way to view the TDE discovery rate.
Figure 16: A stacked histogram of the number of TDEs discovered per year colored by discovery wavelength regime. Telescopes that contributed signficantly to the TDE discovery rate are labelled along the top.

### v2

@@ Appendix A Metadata and References for each TDE Candidate
[274] Forster et al. (2021b) [275] Jones et al. (2021) [276] Munoz-Arancibia et al. (2021a)
[277] Munoz-Arancibia et al. (2021g) [278] Munoz-Arancibia et al. (2021h) [279] Fremling (2021a)
[280] Forster et al. (2021c) [281] Fremling (2021b) [282] Hodgkin et al. (2021)
[283] Wevers et al. (2022) [284] Horesh et al. (2022) [285] Liu et al. (2022)
[286] Perlman et al. (2022) [287] Ravi et al. (2022) [288] Mohan et al. (2022)
[289] Sfaradi et al. (2022) [290] Yao et al. (2022f) [291] Cendes et al. (2022)
[292] Wang et al. (2022) [293] Goodwin et al. (2022) [294] Angus et al. (2022)
[295] Andreoni et al. (2022) [296] Yao et al. (2022a) [297] Fulton et al. (2022b)
[298] Chu et al. (2022) [299] Tucker (2022) [300] Arcavi et al. (2022b)
[301] Arcavi et al. (2022a) [302] Fulton et al. (2022a) [303] Hinkle (2022)
[304] Yao et al. (2022c) [305] Kangas (2022) [306] Moore et al. (2022)
[307] Hammerstein et al. (2022a) [308] Yao et al. (2022e) [309] Somalwar et al. (2022)
[310] Yao et al. (2022b) [311] Blanchard et al. (2022) [312] Hammerstein et al. (2022b)
[313] Fulton et al. (2022c) [314] Perez-Fournon et al. (2022) [315] Yao (2022a)
[316] Yao et al. (2022d) [317] Yao (2022b) [318] Fulton et al. (2022d)
[319] Tonry et al. (2022a) [320] Munoz-Arancibia et al. (2022a) [321] Hodgkin et al. (2022)
[322] Tonry et al. (2022b) [323] Tonry et al. (2022c) [324] De (2022)
[325] Andreoni (2022) [326] Fremling (2022a) [327] Tonry et al. (2022d)
@@
[331] Tonry et al. (2022f) [332] Fremling (2022b) [333] Munoz-Arancibia et al. (2022c)
[334] Stanek (2022) [335] Tonry et al. (2022g) [336] Tonry et al. (2022h)
[337] Fremling (2022c) [338] Fremling (2022d) [339] Fremling (2022e)
[340] Homan et al. (2023) [341] Hammerstein et al. (2023a) [342] Panagiotou et al. (2023)
[343] Zhu et al. (2023) [344] Yao et al. (2023) [345] Hammerstein et al. (2023b)
[346] Aleo et al. (2023) [347] Almeida et al. (2023) [348] Goodwin et al. (2023a)
[349] Rhodes et al. (2023) [350] Goodwin et al. (2023b) [351] Goldtooth et al. (2023)
[352] Yao et al. (2023a) [353] Wise et al. (2023) [354] Chornock et al. (2023)
[355] Hammerstein et al. (2023a) [356] Taguchi et al. (2023b) [357] Yao et al. (2023b)
[358] Taguchi et al. (2023a) [359] Charalampopoulos et al. (2023) [360] Hinkle (2023)
@@
[364] Yao (2023) [365] Chornock (2023) [366] Gomez et al. (2023)
[367] Yao & Qin (2023) [368] Hammerstein et al. (2023b) [369] Stanek (2023)
[370] Tonry et al. (2023a) [371] Tonry et al. (2023b) [372] Godson et al. (2023)
[373] Fremling (2023) [374] Sollerman (2023) [375] Munoz-Arancibia et al. (2023)
[376] Hodgkin et al. (2023) [377] Charalampopoulos et al. (2024) [378] Masterson et al. (2024)
[379] Zhang et al. (2024) [380] Wevers et al. (2024) [381] Huang et al. (2024)
[382] Guolo et al. (2024) [383] Faris et al. (2024b) [384] Cendes et al. (2024)
[385] Dykaar et al. (2024) [386] Christy et al. (2024) [387] Anumarlapudi et al. (2024)
[388] Yao et al. (2024) [389] Newsome et al. (2024) [390] Bykov et al. (2024)
@@
[397] Faris et al. (2024a) [398] Sollerman et al. (2024a) [399] Sollerman et al. (2024b)
[400] Somalwar et al. (2025) [401] Hajela et al. (2025) [402] Yao et al. (2025b)
[403] Sfaradi et al. (2025)
@@ Appendix B Detailed Data Schema
    classification: All reported classifications of this transient stored as a list. Every classification has a confidence associated with it which is derived analytically based on the information available in OTTER. This is described in more detail in §II.4.
@@ B.6 Properties
The classification property is a dictionary with meta-level flag keywords and then all of the classifications stored under the value keyword. The current meta-level flags are spec_classed, unambiguous, and discovery_method, and are discussed in detail in §II.4. In addition to the keywords available in §B.3, classification value elements elements can have: object_class, confidence (which is a flag, as defined in §II.4), and ml_score. An object can have multiple classifications.
@@
Appendix C Non-cumulative TDE candidate Discovery Histogram
Figure 16 shows a non-cumulative version of Figure 7. This is a different way to view the TDE candidate discovery rate.
Figure 16: A stacked histogram of the number of TDE candidates discovered per year colored by discovery wavelength regime. Telescopes that contributed signficantly to the TDE candidate discovery rate are labelled along the top.